# A Spatio-Temporal Machine Learning Model for Mortgage Credit Risk: Default Probabilities and Loan Portfolios

### v2

@@ 3.3 Models considered and implementation details
We consider the following models with increasing levels of complexity: an independent linear hazard model, a linear spatial model, a linear spatio-temporal model, a tree-boosted spatial frailty model, and a tree-boosted spatio-temporal frailty model; see Section 2. We have also tried using an independent linear hazard model with dummy-coded fixed effects for the three-digit postal codes. However, the estimation of this model crashed due to memory overload with the R package stats and numerical instabilities with the R package biglm. For all models that include a spatial or spatio-temporal Gaussian process, we use a Matérn covariance function as defined in Equations (4) and (5) with smoothness parameter $\nu=1.5$ and $m=20$ nearest neighbors for estimation and prediction with Vecchia approximations. We do not estimate the smoothness parameter $\nu$ since “binary data contain no information about [$\nu$]” (De Oliveira, 2000). For purely spatial processes, we use the Euclidean distance to determine the nearest neighbors in Vecchia approximations and a random ordering of the spatial coordinates since this gives accurate approximations (Guinness, 2018). For the spatio-temporal models, we use a correlation-based approach to determine nearest neighbors as in Kang and Katzfuss (2023) and an increasing temporal order and a random spatial order for coordinates with the same time. Following Kang and Katzfuss (2023), the nearest neighbors for the spatio-temporal models are redetermined in every iteration that is a power of two whenever an optimization algorithm is used for determining the parameters $\theta$. For prediction with Vecchia approximations, the observed points appear first in the ordering, and the Gaussian process at prediction points is only conditioned on the training data in the Vecchia approximation. For finding optima for the parameters $\theta$, we use the limited-memory BFGS algorithm. Estimation and prediction with the linear spatial, linear spatio-temporal, tree-boosted spatial, and tree-boosted spatio-temporal frailty models is done using the GPBoost§§§https://github.com/fabsig/GPBoost library version 1.5.8 (Sigrist et al., 2021). The code for reproducing the mortgage credit risk application and generating the data set is publicly available at https://github.com/pkuendig/SpaceTimeML.
@@ 3.5 Prediction of individual default probabilities
We first apply the above-described models to predict one-year-ahead default probabilities of individual mortgage loans for every year from 2008 through 2022. The predictive default probabilities are evaluated using the following prediction accuracy measures: the AUC, the H-measure, the average log-loss, the Brier score, and the expected calibration error (ECE). The AUC measures predictive discrimination ability and can be interpreted as the probability that the predictive probability for a randomly drawn default event is higher than the predictive probability for a randomly drawn loan-year observation without a default. The ECE assesses calibration. A predictive probability $p$ is calibrated if the corresponding event (=default) occurs in $100\times p$ percent. We determine the boundaries of the bins for the ECE by using $20$ equally-spaced quantiles of the empirical distribution of all predictive probabilities of all models and years pooled together. The H-measure, the log-loss, and the Brier score measure overall predictive accuracy. See, e.g., Dimitriadis et al. (2023) for more information on these prediction accuracy measures.
@@
Linear independent 0.7676 0.2024 $7.926\times 10^{-2}$ $1.663\times 10^{-2}$ $1.128\times 10^{-2}$
Linear spatial 0.7702 0.2050 $7.924\times 10^{-2}$ $1.661\times 10^{-2}$ $1.112\times 10^{-2}$
@@
It is important to note that although the observed differences in AUC between the models are relatively small, even small improvements in prediction accuracy yield large economic gains, as the value of the mortgage portfolio of all active loans is in the billions of US dollars. Furthermore, we report in Table 8 in Appendix A.6 p-values of pairwise DeLong tests (DeLong et al., 1988) to assess whether the differences in AUCs of the pooled predictive default probabilities over all test years are significant across the different models. We find that all models have significantly different AUCs at the 5% level, except when comparing the linear independent and linear spatial models.
@@ 3.6 Prediction of loan portfolio loss distributions
[…]
@@ 3.7 Model interpretation
In the following, we aim to better understand the functioning of the tree-boosted and linear spatio-temporal frailty models. In Figure 6, we show the mean of the posterior distribution of the latent spatio-temporal frailty Gaussian process for the tree-boosted spatio-temporal model when training on data up to and including the year 2021, which corresponds to the model for predicting defaults in the most recent year 2022. Due to space constraints, posterior means are not shown for the year 2000. We additionally show these posterior mean maps for the linear spatio-temporal model in Figure 15 in Appendix A.8. We observe that there is considerable variation in the latent frailty variable over space and this variation changes over time. In particular, for the years 2005 and 2006, posterior means for the latent Gaussian variable are high in the New Orleans area, which can be explained by high default rates following Hurricane Katrina in 2005. In 2008 and subsequent years, posterior means are high in the so-called “bubble” states California, Florida, Nevada, and Arizona, where house prices rose particularly rapidly in the run-up to the subprime mortgage crisis (Haughwout et al., 2011). For the year 2020, posterior means are high due to the generally large number of defaults, with the highest values found in the New York City area (not visible at this resolution). This is not surprising given that New York City was one of the first and most affected areas in the U.S. by the COVID-19 pandemic. Comparing posterior means of the tree-boosted spatio-temporal frailty model with those of the linear spatio-temporal model, we find similar patterns over space and time.
@@
For understanding the function $F(\cdot)$ of the tree-boosted frailty models, we use SHAP values and SHAP dependence plots (Lundberg and Lee, 2017). Specifically, we consider SHAP values for the tree-boosted spatio-temporal model trained on data up to the year 2014. We chose this training window because, for the corresponding prediction year 2015, there are pronounced differences in the accuracy of predictive default probabilities between the tree-boosted and linear frailty models. SHAP values are calculated using 10’000 randomly selected instances of the training data. This subsampling is done to reduce the computational complexity and to avoid that SHAP dependence plots are overcrowded. Using a larger random subsample or a different subsample gives almost identical results (results not shown). Figure 8 shows the SHAP values. The predictor variables are ordered according to the average of the absolute values of the SHAP values. According to these results, the eight most important predictor variables in descending order are the credit score (credit_score), the interest rate spread (ir_spread), the current loan-to-value (curr_loan_to_value), the indicator whether multiple borrowers are obligated to repay (multiple_borrowers), the age of the loan (n_months), the indicator whether the mortgage is used for purchase (loan_purpose = P), the original debt-to-income (orig_dept_to_income), and the original unpaid principal balance (orig_unpaid_balance). For the six numeric variables thereof, we report the corresponding SHAP dependence plots in Figure 9. In all dependence plots, we observe strong interaction effects. This can be inferred from the large vertical scatter of the SHAP values and the systematic colored relationship with other predictor variables. For example, the slope of the SHAP values for the credit score variable (credit_score) is less negative if the interest rate spread variable (ir_spread) is positive. I.e., when interest rate spreads are high, the individual credit score matters less compared to when interest rate spreads are low. Furthermore, we find strong non-linear effects for the age of the loan (n_months) and the original unpaid principal balance (orig_unpaid_balance). For instance, the age of a loan is positively related to the default probability up to a certain age of approximately three years, and then the effect flattens out. In addition, the variables interest rate spread (ir_spread), current loan-to-value (curr_loan_to_value), and original debt-to-income (orig_dept_to_income) also have clearly non-linear relationships with slopes that change markedly for large and small values of these variables. For instance, increasing interest rate spreads are related to higher default probabilities, but this effect only holds up to a certain level of approximately two and a half percent, after which the relationship levels off, and additional increases in interest rate spreads are not associated with higher default probabilities. Furthermore, the debt-to-income (orig_dept_to_income) appears to have an approximately logistic-shaped effect on the default probability with the relationship being almost flat for both small debt-to-income values below 20 percent and large values above 50 percent, and in between, the effect of debt-to-income is approximately linear with a slope that interacts, among other things, with the loan age. We conclude that there are strong interaction and non-linear effects. This is likely also the reason why the tree-boosted spatial and spatio-temporal frailty models achieve higher prediction accuracy than their respective linear counterparts.
@@ Acknowledgments
We would like to thank Raffaella Calabrese for her helpful advice for gathering the mortgage data. We also thank Pascal Gantenbein and Christian Kleiber for their useful feedback. This research was partially supported by the Swiss Innovation Agency - Innosuisse (grant number ‘55463.1 IP-ICT’).
@@ A.1 Vecchia approximations for the latent Gaussian process
and 00 otherwise, and the diagonal matrix $D\in\mathbb{R}^{R\times R}$ with $D_{i}$ on the diagonal, we obtain an approximate distribution for the latent Gaussian variable:
@@
The corresponding precision matrix is sparse and given by $\tilde{\Sigma}^{-1}=B^{T}D^{-1}B$. Calculating the Cholesky factor of $\Sigma_{N(r)}$ is the main computational burden of the approximation. Thus, a Vecchia approximation costs $\mathcal{O}(R\times\tilde{m}^{3})$, and saving $B$ requires $O(R\times\tilde{m})$ memory storage. Often, accurate approximations are obtained for small $\tilde{m}$’s.

### v3

@@ 3.3 Models considered and implementation details
We consider the following models with increasing levels of complexity: an independent linear hazard model, a linear spatial model, a linear spatio-temporal model, a tree-boosted spatial frailty model, and a tree-boosted spatio-temporal frailty model; see Section 2. We have also tried using an independent linear hazard model with dummy-coded fixed effects for the three-digit postal codes. However, the estimation of this model crashed due to memory overload with the R package stats and numerical instabilities with the R package biglm. For all models that include a spatial or spatio-temporal Gaussian process, we use a Matérn covariance function as defined in Equations (4) and (5) with smoothness parameter $\nu=1.5$ and $m=20$ nearest neighbors for estimation and prediction with Vecchia approximations. We do not estimate the smoothness parameter $\nu$ since “binary data contain no information about $\nu$]” (de2000bayesian). For purely spatial processes, we use the Euclidean distance to determine the nearest neighbors in Vecchia approximations and a random ordering of the spatial coordinates since this gives accurate approximations (guinness2018permutation). For the spatio-temporal models, we use a correlation-based approach to determine nearest neighbors as in kang2023correlation and an increasing temporal order and a random spatial order for coordinates with the same time. Following kang2023correlation, the nearest neighbors for the spatio-temporal models are redetermined in every iteration that is a power of two whenever an optimization algorithm is used for determining the parameters $\theta$. For prediction with Vecchia approximations, the observed points appear first in the ordering, and the Gaussian process at prediction points is only conditioned on the training data in the Vecchia approximation. For finding optima for the parameters $\theta$, we use the limited-memory BFGS algorithm. Estimation and prediction with the linear spatial, linear spatio-temporal, tree-boosted spatial, and tree-boosted spatio-temporal frailty models is done using the GPBoost§§§https://github.com/fabsig/GPBoost library version 1.5.8 (sigrist2021gpboost). The code for reproducing the mortgage credit risk application and generating the data set is publicly available at https://github.com/pkuendig/SpaceTimeML.
@@ 3.5 Prediction of individual default probabilities
We first apply the above-described models to predict one-year-ahead default probabilities of individual mortgage loans for every year from 2008 through 2022. The predictive default probabilities are evaluated using the following prediction accuracy measures: the AUC, the H-measure, the average log-loss, the Brier score, and the expected calibration error (ECE). The AUC measures predictive discrimination ability and can be interpreted as the probability that the predictive probability for a randomly drawn default event is higher than the predictive probability for a randomly drawn loan-year observation without a default. The ECE assesses calibration. A predictive probability $p$ is calibrated if the corresponding event (=default) occurs in $100\times p$ percent. We determine the boundaries of the bins for the ECE by using $20$ equally-spaced quantiles of the empirical distribution of all predictive probabilities of all models and years pooled together. The H-measure, the log-loss, and the Brier score measure overall predictive accuracy. See, e.g., dimitriadis2023evaluating for more information on these prediction accuracy measures.
@@
Linear independent 0.7676 0.2024 $7.926\times 10^{-2}$ $663\times 10^{-2} 1.1$$128\times 10^{-2}$
Linear spatial 0.7702 0.2050 $7.924\times 10^{-2}$ $661\times 10^{-2} 1.1$$112\times 10^{-2}$
@@
It is important to note that although the observed differences in AUC between the models are relatively small, even small improvements in prediction accuracy yield large economic gains, as the value of the mortgage portfolio of all active loans is in the billions of US dollars. Furthermore, we report in Table 8 in Appendix A.6 p-values of pairwise DeLong tests (delong1988comparing) to assess whether the differences in AUCs of the pooled predictive default probabilities over all test years are significant across the different models. We find that all models have significantly different AUCs at the 5% level, except when comparing the linear independent and linear spatial models.
@@ 3.6 Prediction of loan portfolio loss distributions
The prediction of one-year-ahead loss distributions is performed for every model and year from 2008 to 2022. Quantifying the accuracy of predictive loan portfolio loss distributions is inherently difficult due to the relatively small number of temporally independent observations. Nonetheless, we evaluate the accuracy of the entire predictive loss distributions using the continuous ranked probability score (CRPS) (gneiting2007strictly), of upper 99% predictive quantiles using the 99% quantile loss (koenker1999goodness), and of the mean using the root-mean-square error (RMSE). The CRPS is a proper scoring function that generalizes the absolute error to probabilistic predictions and is defined as $\text{CRPS}(F,\text{L})=\int_{-\infty}^{\infty}(F(y)-\mathds{1}_{\{y\geq\text{L}\}})^{2}\,dy$, where $F$ is a cumulative predictive loan portfolio loss distribution function and $L$ is the realized portfolio loss. Furthermore, upper tails of loss distributions are of particular interest for risk management purposes. We thus consider predictive 99% quantiles and evaluate them using the corresponding quantile loss. The latter is given by $S(q_{\alpha};\text{L})=(\text{L}-q_{\alpha})(\alpha-\mathds{1}_{\{\text{L}\leq q_{\alpha}\}})$, where $q_{\alpha}$ denotes the predictive $\alpha$ quantile. This asymmetric quantile loss function is a proper scoring rule (gneiting2007strictly), and it penalizes observations L which are higher than the predicted quantile $q_{\alpha}$ more heavily. The RMSE is calculated by comparing means of predictive loss distributions with realized portfolio losses.
@@ 3.7 Model interpretation
In the following, we aim to better understand the functioning of the tree-boosted and linear spatio-temporal frailty models. In Figure 6, we show the mean of the posterior distribution of the latent spatio-temporal frailty Gaussian process for the tree-boosted spatio-temporal model when training on data up to and including the year 2021, which corresponds to the model for predicting defaults in the most recent year 2022. Due to space constraints, posterior means are not shown for the year 2000. We additionally show these posterior mean maps for the linear spatio-temporal model in Figure 15 in Appendix A.8. We observe that there is considerable variation in the latent frailty variable over space and this variation changes over time. In particular, for the years 2005 and 2006, posterior means for the latent Gaussian variable are high in the New Orleans area, which can be explained by high default rates following Hurricane Katrina in 2005. In 2008 and subsequent years, posterior means are high in the so-called “bubble” states California, Florida, Nevada, and Arizona, where house prices rose particularly rapidly in the run-up to the subprime mortgage crisis (haughwout2011real). For the year 2020, posterior means are high due to the generally large number of defaults, with the highest values found in the New York City area (not visible at this resolution). This is not surprising given that New York City was one of the first and most affected areas in the U.S. by the COVID-19 pandemic. Comparing posterior means of the tree-boosted spatio-temporal frailty model with those of the linear spatio-temporal model, we find similar patterns over space and time.
@@
For understanding the function $F(\cdot)$ of the tree-boosted frailty models, we use SHAP values and SHAP dependence plots (lundberg2017unified). Specifically, we consider SHAP values for the tree-boosted spatio-temporal model trained on data up to the year 2014. We chose this training window because, for the corresponding prediction year 2015, there are pronounced differences in the accuracy of predictive default probabilities between the tree-boosted and linear frailty models. SHAP values are calculated using 10’000 randomly selected instances of the training data. This subsampling is done to reduce the computational complexity and to avoid that SHAP dependence plots are overcrowded. Using a larger random subsample or a different subsample gives almost identical results (results not shown). Figure 8 shows the SHAP values. The predictor variables are ordered according to the average of the absolute values of the SHAP values. According to these results, the eight most important predictor variables in descending order are the credit score (credit_score), the interest rate spread (ir_spread), the current loan-to-value (curr_loan_to_value), the indicator whether multiple borrowers are obligated to repay (multiple_borrowers), the age of the loan (n_months), the indicator whether the mortgage is used for purchase (loan_purpose = P), the original debt-to-income (orig_dept_to_income), and the original unpaid principal balance (orig_unpaid_balance). For the six numeric variables thereof, we report the corresponding SHAP dependence plots in Figure 9. In all dependence plots, we observe strong interaction effects. This can be inferred from the large vertical scatter of the SHAP values and the systematic colored relationship with other predictor variables. For example, the slope of the SHAP values for the credit score variable (credit_score) is less negative if the interest rate spread variable (ir_spread) is positive. I.e., when interest rate spreads are high, the individual credit score matters less compared to when interest rate spreads are low. Furthermore, we find strong non-linear effects for the age of the loan (n_months) and the original unpaid principal balance (orig_unpaid_balance). For instance, the age of a loan is positively related to the default probability up to a certain age of approximately three years, and then the effect flattens out. In addition, the variables interest rate spread (ir_spread), current loan-to-value (curr_loan_to_value), and original debt-to-income (orig_dept_to_income) also have clearly non-linear relationships with slopes that change markedly for large and small values of these variables. For instance, increasing interest rate spreads are related to higher default probabilities, but this effect only holds up to a certain level of approximately two and a half percent, after which the relationship levels off, and additional increases in interest rate spreads are not associated with higher default probabilities. Furthermore, the debt-to-income (orig_dept_to_income) appears to have an approximately logistic-shaped effect on the default probability with the relationship being almost flat for both small debt-to-income values below 20 percent and large values above 50 percent, and in between, the effect of debt-to-income is approximately linear with a slope that interacts, among other things, with the loan age. We conclude that there are strong interaction and non-linear effects. This is likely also the reason why the tree-boosted spatial and spatio-temporal frailty models achieve higher prediction accuracy than their respective linear counterparts.
@@ Acknowledgments
We would like to thank Raffaella Calabrese for her helpful advice in gathering the mortgage data, Pascal Gantenbein and Christian Kleiber for their useful feedback, and three anonymous reviewers for their constructive comments, which helped to improve this article. This research was partially supported by the Swiss Innovation Agency - Innosuisse (grant number ‘55463.1 IP-ICT’).
@@ A.1 Vecchia approximations for the latent Gaussian process
and 0 otherwise, and the diagonal matrix $D\in\mathbb{R}^{R\times R}$ with $D_{i}$ on the diagonal, we obtain an approximate distribution for the latent Gaussian variable:
$b\overset{\text{approx}}{\sim}\mathcal{N}(0,\tilde{\Sigma}),\penalty 10000\ \penalty 10000\ \penalty 10000\ \tilde{\Sigma}=B^{-1}DB^{-T}.$